OpenStudy (anonymous):

Which of the following ordered pairs is not a solution to the inequality y >= -3x - 2? (0, -1) (0, 0) (-1, 0) (1, 1)

OpenStudy (anonymous):

0, -1

OpenStudy (anonymous):

please check it out again as (-1, 0) is not a solution for this

OpenStudy (anonymous):

Actually its -1, 0. y >= -3x - 2? now put y = -1 and x = 0 -1 >= 0 - 2 -1 >= -2 this is one of the solution because it is true -1 >= -2 but when we put y = 0 and x = -1then 0 >= 3 - 2 0 >= 1 which is not a solution bcuz 0 >= 1 is not true